OpenStudy (anonymous):

Find w(4) for the function w(x) = 3x + 7. A. 41 B. 33 C. 19 D. 14 C?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

you're given : \(\large w(x) = 3x + 7\)

ganeshie8 (ganeshie8):

to find \(\large w(4)\), simply replace \(x \) by \(4\) in given equation

ganeshie8 (ganeshie8):

\(\large w(4) = 3*4 + 7\)

ganeshie8 (ganeshie8):

can u simplify ha ?

OpenStudy (anonymous):

Do i simplify it like an equation?

ganeshie8 (ganeshie8):

nope, just evaluate the right hand side

ganeshie8 (ganeshie8):

\(\large w(4) = 3*4 + 7\) \(\large ~~~~~~~ = 12 + 7\) \(\large ~~~~~~~ = ?\)

OpenStudy (mathmale):

Multiply 3 times 4 FIRST. Once you have that product, add the last constant to it.

OpenStudy (anonymous):

Ok, i did that, i got 19 for the right side

OpenStudy (mathmale):

Why not now find w(0), w(-2) and w(a) for the added practice?

OpenStudy (anonymous):

Wait, so it is C?

OpenStudy (anonymous):

@amistre64

OpenStudy (mathmale):

@blueweek77: Wouldn't it benefit you MUCH more in the long run to learn how to do these problems yourself, and to check them yourself, instead of asking others to check them for you?

ganeshie8 (ganeshie8):

C is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

Thanks @ganeshie8

ganeshie8 (ganeshie8):

np :)